IF,,,,A&B=1100,B&C=1700,C&D=1100,D&E=3300,E&F=5300,F&A=2500,,
THEN..A=?
B=?
C=?
D=?
E=?
F=?

Found 2 solutions by Fombitz, Edwin McCravy:
Answer by Fombitz(32388)

What is the "&" operator stand for?
Please repost your question.

Answer by Edwin McCravy(20060)

  If your "&" means "+", then

A = 2500-k
B = k-1400
C = 3100-k
D = k-2000
E = 5300-k
F = k

You can choose any value for k and 
all the equations check:

A+B=1100, B+C=1700, C+D=1100, 

D+E=3300, E+F=5300, F+A=2500

For example choose k = 2200

A = 2500-2200 = 300
B = 2200-1400 = 800
C = 3100-2200 = 900
D = 2200-2000 = 200
E = 5300-2200 = 3100
F = 2200

Checking:

A+B=1100, B+C=1700, C+D=1100, 

D+E=3300, E+F=5300, F+A=2500

Or choose k = 2350

A = 2500-2350 = 150
B = 2350-1400 = 950
C = 3100-2350 = 750
D = 2350-2000 = 350
E = 5300-2350 = 2950
F = 2350

Checking:

A+B=1100, B+C=1700, C+D=1100, 

D+E=3300, E+F=5300, F+A=2500 

So there are many solutions if "&" is taken to mean "+"

Edwin